# Shape Formation and Locomotion with Joint Movements in the Amoebot Model

### v2

@@ Abstract
In the joint movement extension, the way amoebots move is altered. Amoebots become able to push and pull other amoebots. Feldmann et al. demonstrated the power of joint movements by transforming a line of amoebots into a rhombus within $O(\log n)$ rounds. However, they left the details of the extension open. The goal of this paper is therefore to formalize and extend the joint movement extension. In order to provide a proof of concept for the extension, we consider two fundamental problems of modular robot systems: shape formation and locomotion.
We approach these problems by defining meta-modules of rhombical and hexagonal shape, respectively. The meta-modules are capable of movement primitives like sliding, rotating, and tunneling. This allows us to simulate shape formation algorithms of various modular robot systems. Finally, we construct three amoebot structures capable of locomotion by rolling, crawling, and walking, respectively.
@@
programmable matter, modular robot system, shape formation, locomotion
@@ 1 Introduction
In the reconfigurable circuit extension, the amoebot structure is able to interconnect amoebots by circuits. Each amoebot can send a primitive signal on circuits it is connected to. The signal is received by all amoebots connected to the same circuit. Among others, Feldmann et al. [23] solved the leader election problem, compass alignment problem, and chirality agreement problem within $O(\log n)$ rounds w.h.p. [^1] An event holds with high probability (w.h.p.) if it holds with probability at least $1-1/n^{c}$, where the constant $c$ can be made arbitrarily large. These problems will be important to coordinate the joint movements. Afterwards, Padalkin et al. [45] explored the structural power of the circuits by considering the spanning tree problem and symmetry detection problem. Both problems can be solved within polylogarithmic time w.h.p.
In the joint movement extension, the way amoebots move is altered. In a nutshell, an expanding amoebot is capable of pushing other amoebots away from it, and a contracting amoebot is capable of pulling other amoebots towards it. Feldmann et al. [23] demonstrated the power of joint movements by transforming a line of amoebots into a rhombus within $O(\log n)$ rounds. However, they left the details of the extension open. The goal of this paper is therefore to formalize and extend the joint movement extension. In order to provide a proof of concept for the extension, we consider two fundamental problems of modular robot systems (MRS): shape formation and locomotion. We study these problems from a centralized view to explore the limits of the extension.
In the shape formation problem, an MRS has to reconfigure its structure into a given shape. Examples for shape formation algorithms in the original amoebot model can be found in [17, 18, 35, 40]. However, all of these are subject of the aforementioned natural lower bound. To our knowledge, polylogarithmic time solutions were found for two types of MRSs: in the nubot model [60] and crystalline atom model [6]. We will show that in the joint movement extension, the amoebots are able to simulate the shape formation algorithm for the crystalline atom model, and with that break the lower bound.
@@ 2.1 Geometric Amoebot Model
In the geometric amoebot model [15], a set of $n$ amoebots is placed on the infinite regular triangular grid graph $G_{\Delta}=(V,E)$ (see Figure 0(a)). An amoebot is an anonymous, randomized finite state machine that either occupies one or two adjacent nodes of $G_{\Delta}$, and every node of $G_{\Delta}$ is occupied by at most one amoebot. If an amoebot occupies just one node, it is called contracted (a single node) and otherwise expanded (two nodes connected by a black edge). Two amoebots that occupy adjacent nodes in $G_{\Delta}$ are connected by bonds (red edges). An amoebot can move through contractions and expansions (see Figure 2). We refer to [15] for more details.
(a) Triangular grid.
(b) Rhombical tesselation.
(c) Hexagonal tesselation.
Figure 1: Domains. The figures show the domains we are working with. The tesselations will be explained in Section 3. Note that there are spaces between the meta-modules since a node cannot be occupied by more than one amoebot. However, the spaces do not contain any nodes.
@@
Figure 2: Movements in the classical model. Red lines indicate bonds. Blue amoebots are expanding. Green amoebots are contracting. (a) An expanded amoebot occupying nodes $v$ and $w$ can contract into one of its occupied nodes. Thereafter, it only occupies that node. (b) A contracted amoebot occupying node $u$ can expand into an unoccupied adjacent node $v$. Thereafter, it occupies both, $u$ and $v$. (c) Two adjacent amoebots can perform a handover if one is contracted and the other is expanded: let node $u$ be occupied by the contracted amoebot, and let nodes $v$ and $w$ be occupied by the expanded amoebot such that nodes $u$ and $v$ are adjacent. The contracted amoebot expands into node $v$ while the expanded node contracts into node $w$.
Let the amoebot structure $S\subseteq V$ be the set of nodes occupied by the amoebots. We say that $S$ is connected iff $G_{S}$ is connected, where $G_{S}=G_{\Delta}|_{S}$ is the graph induced by $S$. Initially, $S$ is connected. Also, we assume the fully synchronous activation model, i.e., the time is divided into synchronous rounds, and every amoebot is active in each round. We justify this assumption with the reconfigurable circuit extension. On activation, each amoebot may perform a movement and update its state as a function of its previous state. However, if an amoebot fails to perform its movement, it remains in its previous state. The time complexity of an algorithm is measured by the number of synchronized rounds required by it.
@@ 2.2 Joint Movement Extension
In the joint movement extension [23], the way the amoebots move is altered. The idea behind the extension is to allow amoebots to push and pull other amoebots. The necessary coordination of such movements can be provided by the reconfigurable circuit extension. Since we study the problems from a centralized view, we refrain from describing that extension and refer to [23, 45]. In the following, we formalize the joint movement extension. Joint movements are performed in two steps.
@@
In the first step, the amoebots remove bonds from $G_{S}$ as follows. Each amoebot can decide to release an arbitrary subset of its currently incident bonds in $G_{S}$. A bond is removed iff either of the amoebots at the endpoints releases the bond. However, an expanded amoebot cannot release the bond connecting its occupied nodes. Let $E_{R}\subseteq E_{S}$ denote the set of the remaining bonds, and $G_{R}=(S,E_{R})$ the resulting graph. We require that $G_{R}$ is connected since otherwise, disconnected parts might float apart. We say that a connectivity conflict occurs iff $G_{R}$ is not connected. Whenever a connectivity conflict occurs, the amoebot structure transitions into an undefined state such that we become unable to make any statements about the structure.
In the second step, each amoebot may perform one of the following movements within the time period $[0,1]$. A contracted amoebot may expand on one of the axes as follows (see Figure 2(a)). At $t=0$, the amoebot splits its single node into two nodes, adds a bond between them, and assigns each of its other incident bonds to one of the resulting nodes. Then, the amoebot uniformly extends the length of the new bond. At $t\in[0,1]$, that bond has a length of $t$. In the process, the incident bonds do not change their orientations or lengths.
@@
Figure 3: Movements in the extension. Red lines indicate bonds. Blue amoebots are expanding. Green amoebots are contracting. The first two figures show a movement in $0.5$ time steps.
An expanded amoebot may contract into a single node as follows (see Figure 2(b)). The amoebot uniformly reduces the length of the bond between its endpoints. At $t\in[0,1]$, the bond has a length of $1-t$. At $t=1$, the amoebot removes that bond, fuses its endpoints into a single node, and assigns all of its other incident bonds to the resulting node. In the process, the incident bonds do not change their orientations or lengths. We remove duplicates of bonds if any occur (see Figure 3(a)).
(a) Duplicates.
(b) Collision.
Figure 4: Special cases. Red lines indicate bonds. Blue amoebots are expanding. Green amoebots are contracting. Each figure shows a movement in $0.5$ time steps.
Furthermore, a contracted amoebot $x$ occupying node $u$ and an expanded amoebot $y$ occupying nodes $v$ and $w$ may perform a handover if there is a bond $b$ between $u$ and $v$, as follows (see Figure 2(c)). At an arbitrary $t\in[0,1]$, we flip $b=\{u,v\}$ and $\{v,w\}$ such that $x$ becomes an expanded amoebot occupying nodes $u$ and $v$, $y$ becomes a contracted amoebot occupying $w$, and $b$ becomes $\{v,w\}$. We have to include the handover to ensure universality of the model since otherwise, it would not be possible to move through a narrow tunnel. In theory, all movement primitives presented in Section 3 can be realized without handovers. However, for reasons of clarity and comprehensibility, we will still make use of the handovers.
@@
The amoebots may not be able to perform their movements. We distinguish between two cases. First, the amoebots may not be able to perform their movements while maintaining their relative positions (see Figure 4(a)). We call that a structural conflict. Second, parts of the structure may collide into each other. More precisely, a collision occurs if there is a $t\in[0,1]$ such that two non-adjacent bonds intersect at some point (see Figure 4(b)). In particular, note that an expanded amoebot cannot contract if it has two incident bonds to the same node even if the amoebot occupying that node expands at the same time without causing a collision (see Figure 3(b)). Whenever either a structural conflict or a collision occurs, the amoebot structure transitions into an undefined state such that we become unable to make any statements about the structure.
Otherwise, at $t=1$, we obtain a graph $G_{M}=(S,E_{M})$ that can be mapped on the triangular grid $G_{\Delta}$ (see Figure 4(a)). Since $G_{R}$ is connected, the mapping of $G_{M}$ is unique except for translations (if one exists). We choose any mapping as our next amoebot structure. Afterwards, the amoebots reestablish all possible bonds. Unless stated otherwise, each arrow in the figures indicates a single round. The left side shows the structure after the removal of bonds (first step), and the right side the structure after the execution of movements (second step). We chain multiple rounds if we do not change bonds.
(a) Structural conflicts and mappings. In the left figure, the expansions cause a structural conflict. In the right figure, the amoebot structure can be mapped onto the triangular grid after the movement.
(b) Collision. Initially, we have a valid amoebot structure given ($t=0$). The resulting amoebot structure can be mapped on $G_{\Delta}$ ($t=1$). However, for $t\in[0.25,0.75]$, parts of the structure collide.
Figure 5: Joint movements. Red lines indicate bonds. Blue amoebots are expanding.
We assume that the joint movements are performed within look-compute-move cycles. In the look phase, each amoebot observes its neighborhood and receives signals (beeps) from other amoebots, according to the reconfigurable circuit structure of the system. In the compute phase, each amoebot may perform computations, change its state, and decide the actions to perform in the next phase (i.e., which bonds to release, and which movement to perform). In the move phase, each amoebot may release an arbitrary subset of its incident bonds, and perform a movement.

### v3

# Reconfiguration and Locomotion with Joint Movements in the Amoebot Model

@@ Abstract
In the joint movement extension, the way amoebots move is altered. Amoebots become able to push and pull other amoebots. Feldmann et al. demonstrated the power of joint movements by transforming a line of amoebots into a rhombus within $O(\log n)$ rounds. However, they left the details of the extension open. The goal of this paper is therefore to formalize and extend the joint movement extension. In order to provide a proof of concept for the extension, we consider two fundamental problems of modular robot systems: reconfiguration and locomotion.
We approach these problems by defining meta-modules of rhombical and hexagonal shape, respectively. The meta-modules are capable of movement primitives like sliding, rotating, and tunneling. This allows us to simulate reconfiguration algorithms of various modular robot systems. Finally, we construct three amoebot structures capable of locomotion by rolling, crawling, and walking, respectively.
@@
amoebot model, programmable matter, modular robot system, reconfiguration, locomotion
@@ 1 Introduction
In the reconfigurable circuit extension, the amoebot structure is able to interconnect amoebots by circuits. Each amoebot can send a primitive signal on circuits it is connected to. The signal is received by all amoebots connected to the same circuit. Among others, Feldmann et al. [23] solved the leader election problem, compass alignment problem, and chirality agreement problem within $O(\log n)$ rounds w.h.p. [^1] An event holds with high probability (w.h.p.) if it holds with probability at least $1-1/n^{c}$, where the constant $c$ can be made arbitrarily large. These problems will be important to coordinate the joint movements. Afterwards, Padalkin et al. [46, 47] proposed polylogarithmic time algorithms to various problems including spanning tree computation, symmetry detection, and shortest path computation.
In the joint movement extension, the way amoebots move is altered. In a nutshell, an expanding amoebot is capable of pushing other amoebots away from it, and a contracting amoebot is capable of pulling other amoebots towards it. Feldmann et al. [23] demonstrated the power of joint movements by transforming a line of amoebots into a rhombus within $O(\log n)$ rounds. However, they left the details of the extension open. The goal of this paper is therefore to formalize and extend the joint movement extension. In order to provide a proof of concept for the extension, we consider two fundamental problems of modular robot systems (MRS): reconfiguration and locomotion. We study these problems from a centralized view to explore the limits of the extension.
In the reconfiguration problem, an MRS has to reconfigure its structure into a given shape. Examples for reconfiguration algorithms in the original amoebot model can be found in [17, 18, 36, 41]. However, all of these are subject of the aforementioned natural lower bound. To our knowledge, polylogarithmic time solutions were found for two types of MRSs: in the nubot model [62] and crystalline atom model [6]. We will show that in the joint movement extension, the amoebots are able to simulate the reconfiguration algorithm for the crystalline atom model, and with that break the lower bound of the original amoebot model.
@@ 2.1 Geometric Amoebot Model
In this section, we introduce the geometric amoebot model [15]. We slightly deviate from the original model to make it more suitable to our extension. A set of $n$ amoebots is placed on the infinite regular triangular grid graph $G_{\Delta}=(V_{\Delta},E_{\Delta})$ (see Figure 1). An amoebot is an anonymous, randomized finite state machine in the form of a line segment. The endpoints may either occupy the same node or two adjacent nodes. If the endpoints occupy the same node, the amoebot has length 00 and is called contracted and otherwise, it has length $1$ and is called expanded. Every node of $G_{\Delta}$ is occupied by at most one amoebot. Two endpoints of different amoebots that occupy adjacent nodes in $G_{\Delta}$ are connected by bonds (red edges). Let $V\subseteq V_{\Delta}$ be the nodes occupied by the amoebots. We define graph $S=G_{\Delta}|_{V}$ induced by $V$ to be the connectivity graph of the amoebot structure. Initially, $S$ is connected and we require that $S$ stays connected at all times.
Figure 1: Amoebot structure. The amoebots are shown in black. Red edges indicate bonds.
An amoebot can move through expansions and contractions. A contracted amoebot occupying node $u$ can expand into an unoccupied adjacent node $v$ (see the blue amoebot in Figure 2(a)). Thereafter, it occupies both, $u$ and $v$. An expanded amoebot occupying nodes $v$ and $w$ can contract into one of its occupied nodes (see the green amoebot in Figure 2(b)). Thereafter, it only occupies that node. Furthermore, let a contracted amoebot $x$ occupy node $u$ and an expanded amoebot $y$ occupy nodes $v$ and $w$ such that $u$ and $v$ are adjacent. These amoebots can perform a handover where $x$ expands into $v$ and $y$ contracts into $w$ simultaneously (see Figure 2(c) where $x$ is marked in blue and $y$ is marked in green). Thereafter, $x$ occupies $u$ and $v$, and $y$ occupies $w$.
We refer to [15] for more details.
@@
Figure 2: Movements in the classical model. Red edges indicate bonds. Blue amoebots are expanding. Green amoebots are contracting.
@@ 2.2 Joint Movement Extension
In the joint movement extension [23], the way the amoebots move is altered. The idea behind the extension is to allow amoebots to push and pull other amoebots. In the following, we formalize the joint movement extension.
We first redefine the connectivity graph of an amoebot structure. In the graph, we represent each amoebot and bond by an edge. Let $A$ denote the set of all edges representing amoebots, $V$ the set of all their endpoints, and $B$ the set of edges representing bonds between these endpoints. The edges of $A$ are pairwise disjoint. We define the connectivity graph of an amoebot structure as the graph $S=(V,A\cup B)$.
Note that the connectivity graph does not contain any information about the positions of the amoebots in the amoebot structure. Each amoebot contributes one edge to $A$ and two nodes to $V$ even if it is in a contracted state.
To each edge in $A\cup B$, we assign a geometric orientation and a length, such that we obtain line segments. We only allow orientations in parallel to the axes of the triangular grid $G_{\Delta}$. Each amoebot has a length in $[0,1]$ and each bond has a length of $1$. We consider two edges in $A\cup B$ as adjacent if and only if they share a common endpoint or they are connected by another edge of length 00 (bonds connected by a contracted amoebot). The amoebot structure is defined by its connectivity graph and the assignment of orientations and lengths. For the sake of simplicity, we will denote each amoebot structure by its connectivity graph.
We call an amoebot structure connected if and only if $S$ is connected. We call an amoebot structure valid if and only if we can embed the graph in the plane $\mathbb{R}^{2}$ in compliance with the orientations and lengths of all amoebots and bonds (i.e., $A\cup B$) such that only adjacent edges intersect. Note that an embedding is unique except for translations if and only if it is connected. We call a valid amoebot structure triangular if and only if the embedding can be aligned with the triangular grid. Note that this is the case if and only if each amoebot in $A$ has either length 00 or $1$. We call a valid amoebot structure complete if and only if there is no other valid amoebot structure $S^{\prime}=(V,A\cup B^{\prime})$ with $B\subsetneq B^{\prime}$. We assume that initially, the amoebot structure $S_{0}=(V,A\cup B_{0})$ is connected, valid, triangular, and complete.
We assume the fully synchronous activation model, i.e., the time is divided into synchronous rounds, and every amoebot is active in each round. Furthermore, we make the idealistic assumption that all movements start at the same time and are performed at the same speed. W.l.o.g., we assume that all movements occur within the time period $[0,1]$.
Let $S_{i}=(V,A\cup B_{i})$ be an connected, valid, triangular, and complete amoebot structure. Joint movements are performed in four steps. In the first step, the amoebots remove bonds from $S_{i}$ as follows. Each amoebot can decide to release an arbitrary subset of its currently incident bonds in $S_{i}$. A bond is removed if and only if either of the amoebots at the endpoints releases the bond. Note that we only remove bonds (i.e., edges in $B_{i}$) but no amoebots (i.e., edges in $A$). Let $R_{i}\subseteq B_{i}$ be the set of the remaining bonds and $S^{0}_{i}=(V,A\cup R_{i})$ the resulting amoebot structure.
@@
We require that $S^{0}_{i}$ is connected since otherwise, disconnected parts might float apart. We say that a connectivity conflict occurs if and only if $S^{0}_{i}$ is not connected. Whenever a connectivity conflict occurs, the amoebot structure transitions into an undefined state such that we become unable to make any statements about the structure. Note that $S^{0}_{i}$ is still valid and triangular since $S^{0}_{i}$ is a subgraph of $S_{i}$. However, $S^{0}_{i}$ is not complete anymore if at least one bond was removed.
In the second step, each amoebot may perform one of the following movements. A contracted amoebot may expand on one of the axes as follows (see blue amoebot in Figure 3(a)). At $t=0$, the amoebot can reorientate itself and reassign each of its incident bonds to one of its endpoints. Bonds assigned to an endpoint will stay with that endpoint as the amoebot expands. At $t\in[0,1]$, the amoebots has a length of $t$. In the process, the incident bonds do not change their orientations or lengths. As a result, the expanding amoebot pushes all connected amoebots. An expanded amoebot may contract analogously by reversing the contraction (see green amoebot in Figure 3(b)). Thereby, it pulls all connected amoebots.
@@
Figure 3: Movements in the extension. Red edges indicate bonds. Blue amoebots are expanding. Green amoebots are contracting. The figures show the movements in $0.5$ time steps.
@@
Let $S^{t}_{i}=(V,A\cup R_{i})$ be the amoebot structure at $t\in[0,1]$. Note that for all $t\in[0,1]$, $S^{t}_{i}$ is connected since $S^{0}_{i}$ is connected. We require that $S^{t}_{i}$ is valid for all $t\in[0,1]$. If there is a $t\in(0,1]$ such that $S^{t}_{i}$ is not valid, the amoebot structure transitions into an undefined state such that we become unable to make any statements about the structure. We distinguish between two cases. First, we say that a structural conflict occurs if and only if we cannot embed $S^{t}_{i}$ in the plane, i.e., the amoebots are not able to maintain their relative positions (see Figure 4). Second, we say that a collision occurs if and only if we can embed $S^{t}_{i}$ in the plane, but there is a forbidden intersection, i.e., two non-adjacent edges intersect or two adjacent edges intersect at a point which is not their endpoints (see Figure 5).
Figure 4: Structural conflict. We are not able to expand the blue amoebots horizontally without tearing up the amoebot structure. Hence, the expansions cause a structural conflict and the amoebot structure transitions into an undefined state.
Figure 5: Collision. Initially, we have a valid amoebot structure given ($t=0$). We expand the blue amoebots horizontally. Note that the amoebot structure is also valid at $t=1$. However, for $t\in[0.25,0.75]$, parts of the amoebot structure collide. Hence, the amoebot structure transitions into an undefined state.
The detection of structural conflicts and collisions is not within the scope of this paper simply because we only consider movements where structural conflicts and collisions cannot occur. We refer to [26] for more details.
Suppose that $S^{t}_{i}$ is valid for all $t\in[0,1]$. Note that for all $t\in(0,1)$, $S^{t}_{i}$ is no longer triangular if there is at least one expansion or contraction. At $t=1$, we obtain an amoebot structure $S^{1}_{i}$ which is triangular since each amoebot in $A$ has length 00 or $1$ again. However, $S^{1}_{i}$ is not necessarily complete.
Each arrow in the figures of Sections 3, 4, 5 and 6 indicates the execution of expansions and contractions. The left side shows $S^{0}_{i}$ (i.e., the amoebot structure after the removal of bonds in the first step), and the right side $S^{1}_{i}$ (i.e., the amoebot structure after the execution of movements in the second step).
In the third step, the amoebots reestablish bonds until the amoebot structure becomes complete again. Let $B^{\prime}_{i}$ be the set of bonds afterwards. We obtain the amoebot structure $S^{\prime}_{i}=(V,A\cup B^{\prime}_{i})$. Note that $S^{\prime}_{i}$ is connected, valid, triangular, and complete.
In the last step, pairs of amoebots may perform isolated handovers as follows. Note that each amoebot can participate in only one handover.
Consider a contracted amoebot $x$ with endpoints $x_{1}$ and $x_{2}$ and an expanded amoebot $y$ with endpoints $y_{1}$ and $y_{2}$, which are connected by a bond $b$ with endpoints $x_{2}$ and $y_{2}$ (see Figure 6 where $x$ is marked in blue and $y$ is marked in green). Intuitively, we want to switch the association of the expanded amoebot and the bond. More precisely, (i) amoebot $x$ becomes expanded (i.e., it has length $1$) and it changes its orientation to the one of the bond $b$, (ii) amoebot $y$ becomes contracted (i.e., it has length 00), (iii) bond $b$ changes its orientation to the one of the previously expanded amoebot, (iv) all bond previously adjacent to $x_{2}$ become adjacent to $x_{1}$, and (v) all bonds previously adjacent to $y_{2}$ except for $b$ become adjacent to $x_{2}$.
Let $B_{i+1}$ be the set of bonds after the handovers and $S_{i+1}=(V,A\cup B_{i+1})$ the resulting amoebot structure. By construction of the handovers, $S_{i+1}$ has the same properties as $S^{\prime}_{i}$, i.e., it is connected, valid, triangular, and complete. We call a movement valid if and only if we are able to perform all four steps, i.e., the amoebot structure did not transition into an undefined state.
Figure 6: Handover. Red edges indicate bonds. The blue amoebot becomes expanded. The green amoebot becomes contracted.
Remark 2.1.
We have to include the handover to ensure compatibility with algorithms for the original model and universality of the model since otherwise, it would not be possible to move through a narrow tunnel. However, it is not possible to define the handover as a movement in the second step without breaking previous assumptions.
If both amoebots performed their movements simultaneously, the bond between them would change its orientation and length during the movement (see Figure 7(a) where $x$ is marked in blue and $y$ is marked in green). Also, note that we cannot prevent the contracting amoebot from pulling all other incident amoebots with it such that the handover would not be an isolated local operation. This may result in structural conflicts or collisions.
If the amoebots performed their movements sequentially, they would have to move faster than amoebots performing an independent movement, and they would intersect during the movement (see Figure 7(b) where $x$ is marked in blue and $y$ is marked in green).
Since all these definitions of handovers are breaking our previous assumptions, we have decided to not include handovers as a movement in the second step and include an extra step for them. We will refrain from using them in all of our following constructions.
(a) Simultaneous handover. The second shows a handover in $0.5$ time steps.
(b) Sequential handover. The figure shows a handover in $0.25$ time steps.
Figure 7: Possible implementations of a handover as a movement in the second step. Red edges indicate bonds. Blue amoebots are expanding. Green amoebots are contracting.
We call a sequence of valid movements a reconfiguration. For the sake of simplicity, we will denote each reconfiguration by the sequence of the amoebot structures between the movements, i.e., $(S_{0},S_{1},\dots,S_{F})$ denotes a reconfiguration from $S_{0}$ to $S_{F}$. Note that we hide the exact movements by doing so.
{observation}
Each valid movement is reversible, i.e., if there is a valid movement from $S_{i}$ to $S_{i+1}$, then there is also a valid movement from $S_{i+1}$ to $S_{i}$. In order to obtain the latter from the former, we proceed as follows. In the first step, we remove the bonds that we have added previously in the third round. In the second step, each amoebot performs the reverse movement. In the third step, we add the bonds that we have removed previously in the first round. We execute the fourth step before the first step where the amoebots perform the reverse handovers.
The reversibility of valid movements immediately implies that each reconfiguration is reversible, i.e., if there is a reconfiguration $(S_{0},\dots,S_{F})$, then there is also a reconfiguration $(S_{F},\dots,S_{0})$.
In this paper, we assume that we have a centralized scheduler. The scheduler knows the current state of the amoebot structure at all times. At the beginning of each synchronous round, it decides for each amoebot (i) which bonds to release, (ii) which movements to perform, (iii) which bonds to reestablished and (iv) which handovers to perform. We leave the design of distributed solutions for future work.